Suppose you deposited $11,000 in a savings account earning 3.6% interest compounding daily. How long will it take for the balance to grow to $22,000? Answer in years rounded to one decimal place.
Answer #1

We use the formula:
A=P(1+r/365)^365n
where
A=future value
P=present value
r=rate of interest
n=time period.

22000=11000*(1+0.036/365)^365n

(22000/11000)=(1+0.036/365)^365n

2=(1.00009863)^365n

Taking log on both sides;

log 2=365n*log 1.00009863

n=1/365[log 2/log 1.00009863]

=19.3 years(Approx).